# On the regularity of axisymmetric, swirl-free solutions of the Euler equation in four and higher dimensions

### v3

@@ 1 Introduction
As further evidence of potential singularity formation in very high dimensions, we prove that any solution of the Euler equation in $\mathbb{R}^{d}$ with a particular geometry—axisymmetric, swirl-free, with a scalar vorticity $\omega^{0}(r,z)$ that is odd in $z$ and satisfying $\omega^{0}(r,z)\geq 0$ for all r,z>0r,z>0— blows up in finite time as long as a large enough proportion of energy remains in the vertical velocity component. Furthermore, the proportion of energy remaining in the vertical component necessary to guarantee finite-time blowup shrinks at a rate $\frac{1}{d}$, making the requirement very weak in high dimensions.
@@ 3.1 Preliminaries
$\nabla u(x)=\partial_{r}u_{r}(r,z)e_{r}\otimes e_{r}+\partial_{z}u_{z}(r,z)e_{z}\otimes e_{z}+\partial_{r}u_{z}(r,z)e_{r}\otimes e_{z}+\partial_{z}u_{r}(r,z)e_{z}\otimes e_{r}+u_{r}(r,z)\nabla e_{r}.$ (3.5)
@@
$\Delta u=-\partial_{z}^{2}u_{r}e_{r}-\partial_{z}^{2}u_{z}e_{z}-\partial_{r}^{2}u_{z}e_{z}-\frac{k}{r}\partial_{r}u_{z}e_{z}-\operatorname{div}\left(\partial_{r}u_{r}e_{r}\otimes e_{r}\right)-\operatorname{div}\left(\frac{u_{r}}{r}\tilde{I}_{d}\right).$ (3.30)
@@
It is obvious that $\nabla\nabla\cdot(-\Delta)^{-1}v\in H^{s}_{as\&gr}$, so it remains only to show that $2\operatorname{div}\nabla_{asym}(-\Delta)^{-1}v\in H^{s}_{*}$. Letting $A=\nabla_{asym}(-\Delta)^{-1}v$, we can see that
@@
We note here that the above proposition shows that for all $v\in H^{s}_{as},P_{df}(v)\in H^{s}_{as}.$
@@ 3.2 Local existence of solutions
and then by proving strong convergence $u^{\epsilon}\to u$ in $C\left([0,T_{max}),H^{s}_{df}\left(\mathbb{R}^{d}\right)\right)\cap C^{1}\left([0,T_{max}),H^{s-1}_{df}\left(\mathbb{R}^{d}\right)\right)$. Note that $\mathcal{J}_{\epsilon}$ is a regularization operator given by convolution with a smooth, positive, compactly supported function of mass $1$. As long as $\mathcal{J}_{\epsilon}$ involves convolution with a radial function, then it will preserve the class of axisymmetric, swirl-free vector fields. The Picard iteration then guarantees that $u^{\epsilon}$ must be axisymmetric and swirl-free. Finally, $u$ must be axisymmetric and swirl-free, because it is trivial that this property is preserved by strong convergence in $H^{s}$.
@@ 4 The Biot-Savart law in higher dimensions
$\int_{\mathbb{S}^{d-2}}f(y_{1})\mathop{}\!\mathrm{d}S(y)=m_{d-3}\int_{-1}^{1}\left(1-y_{1}^{2}\right)^{\frac{d-4}{2}}f(y_{1})\mathop{}\!\mathrm{d}y_{1}=m_{d-3}\int_{0}^{\pi}f(\cos\theta)\sin^{d-3}\theta\mathop{}\!\mathrm{d}\theta,$ (4.2)
@@
$\int_{\mathbb{S}^{d-2}}f(y_{1})\mathop{}\!\mathrm{d}S(y)=\int_{-1}^{1}\int_{|\tilde{y}|=\sqrt{1-y_{1}^{2}}}f(y_{1})\frac{1}{\sqrt{1-y_{1}^{2}}}\mathop{}\!\mathrm{d}S(\tilde{y})\mathop{}\!\mathrm{d}y_{1}=\int_{-1}^{1}f(y_{1})\frac{1}{\sqrt{1-y_{1}^{2}}}\left|\mathbb{S}^{d-3}_{\sqrt{1-y_{1}^{2}}}\right|\mathop{}\!\mathrm{d}y_{1}.$
@@
$u_{r}(r,z)=-\frac{d-2}{2\pi}\int_{-\infty}^{\infty}\int_{0}^{\infty}\bar{r}^{d-2}(z-\bar{z})\omega(\bar{r},\bar{z})\int_{-1}^{1}\frac{y_{1}(1-y_{1}^{2})^{\frac{d-4}{2}}}{\left(r^{2}+\bar{r}^{2}-2r\bar{r}y_{1}+(z-\bar{z})^{2}\right)^{\frac{d}{2}}}\mathop{}\!\mathrm{d}y_{1}\mathop{}\!\mathrm{d}\bar{r}\mathop{}\!\mathrm{d}\bar{z}$ (4.3)
@@
$u_{r}(r,z)=-\frac{d-2}{2\pi}\int_{-\infty}^{\infty}\int_{0}^{\infty}\bar{r}^{d-2}(z-\bar{z})\omega(\bar{r},\bar{z})\int_{0}^{\pi}\frac{\sin^{d-3}\theta\cos\theta}{\left(r^{2}+\bar{r}^{2}-2r\bar{r}\cos\theta+(z-\bar{z})^{2}\right)^{\frac{d}{2}}}\mathop{}\!\mathrm{d}\theta\mathop{}\!\mathrm{d}\bar{r}\mathop{}\!\mathrm{d}\bar{z}$ (4.4)
@@
$u_{z}(r,z)=\frac{d-2}{2\pi}\int_{-\infty}^{\infty}\int_{0}^{\infty}\bar{r}^{d-2}\omega(\bar{r},\bar{z})\int_{0}^{\pi}\frac{\sin^{d-3}\theta(r\cos\theta-\bar{r})}{\left(r^{2}+\bar{r}^{2}-2r\bar{r}\cos\theta+(z-\bar{z})^{2}\right)^{\frac{d}{2}}}\mathop{}\!\mathrm{d}\theta\mathop{}\!\mathrm{d}\bar{r}\mathop{}\!\mathrm{d}\bar{z}.$ (4.6)
@@
$u_{r}(r,z)=-\alpha_{d}\int_{-\infty}^{\infty}\int_{0}^{\infty}\bar{r}^{d-2}(z-\bar{z})\omega(\bar{r},\bar{z})\int_{\mathbb{S}^{d-2}}\frac{y_{1}}{\left(r^{2}+\bar{r}^{2}-2r\bar{r}y_{1}+(z-\bar{z})^{2}\right)^{\frac{d}{2}}}\mathop{}\!\mathrm{d}S(y)\mathop{}\!\mathrm{d}\bar{r}\mathop{}\!\mathrm{d}\bar{z}.$
@@ 5 Conditional regularity in four and higher dimensions
$\|\omega\circ X(\cdot,t)\|_{L^{q}\left(\mathbb{R}^{d}\right)}^{q}$ $=\|\omega\circ X(\cdot,t)\|_{L^{q}\left(\mathcal{C}_{R}\right)}^{q}+\|\omega\circ X(\cdot,t)\|_{L^{q}\left(\mathcal{C}_{R}^{c}\right)}^{q}$ (5.20)
$\leq\left(\left\|\frac{\omega^{0}}{r^{k}}\right\|_{L^{q}\left(\mathcal{C}_{R}\right)}^{q}R^{kq}+\left\|\omega^{0}\right\|_{L^{q}\left(\mathcal{C}_{R}^{c}\right)}^{q}\right)\left(1+\frac{1}{R}\int_{0}^{t}\|u_{r}^{+}(\cdot,\tau)\|_{L^{\infty}}\mathop{}\!\mathrm{d}\tau\right)^{kq}$ (5.21)
@@
$\|u_{r}^{+}(\cdot,t)\|_{L^{\infty}}\leq C_{d}\left\|\frac{\omega^{0}}{r^{k}}\right\|_{L^{\infty}}^{\frac{1}{2}}\left(\left\|\omega^{0}\right\|_{L^{1}(\mathcal{C}_{R}^{c})}+R^{k}\left\|\frac{\omega^{0}}{r^{k}}\right\|_{L^{1}(\mathcal{C}_{R})}\right)^{\frac{1}{2}}\left(1+\frac{1}{R}\int_{0}^{t}\|u_{r}^{+}(\cdot,\tau)\|_{L^{\infty}}\mathop{}\!\mathrm{d}\tau\right)^{\frac{k}{2}}.$ (5.23)
@@
$\|u_{r}^{+}(\cdot,t)\|_{L^{\infty}}$ $\leq C_{d}\left\|\frac{\omega(\cdot,t)}{r^{k}}\right\|_{L^{\infty}}^{\frac{1}{2}}\|\omega(\cdot,t)\|_{L^{1}}^{\frac{1}{2}}$ (5.27)
$\leq\left\|\frac{\omega^{0}}{r^{k}}\right\|_{L^{\infty}}^{\frac{1}{2}}\left(\left\|\omega^{0}\right\|_{L^{1}(\mathcal{C}_{R}^{c})}+R^{k}\left\|\frac{\omega^{0}}{r^{k}}\right\|_{L^{1}(\mathcal{C}_{R})}\right)^{\frac{1}{2}}\left(1+\frac{1}{R}\int_{0}^{t}\|u_{r}^{+}(\cdot,\tau)\|_{L^{\infty}}\mathop{}\!\mathrm{d}\tau\right)^{\frac{k}{2}}.$ (5.28)
@@ 6 Anti-parallel vortex tubes
$H(r,z,\bar{r},\bar{z})=\int_{0}^{1}\tau\left(1-\tau^{2}\right)^{\frac{d-4}{2}}\Bigg(\frac{z+\bar{z}}{\left(r^{2}+\bar{r}^{2}-2r\bar{r}\tau+(z+\bar{z})^{2}\right)^{\frac{d}{2}}}+\frac{z-\bar{z}}{\left(r^{2}+\bar{r}^{2}+2r\bar{r}\tau+(z-\bar{z})^{2}\right)^{\frac{d}{2}}}\\ -\frac{z+\bar{z}}{\left(r^{2}+\bar{r}^{2}+2r\bar{r}\tau+(z+\bar{z})^{2}\right)^{\frac{d}{2}}}-\frac{z-\bar{z}}{\left(r^{2}+\bar{r}^{2}-2r\bar{r}\tau+(z-\bar{z})^{2}\right)^{\frac{d}{2}}}\Bigg)\mathop{}\!\mathrm{d}\tau,$ (6.3)
@@
$G(r,z,\bar{r},\bar{z})=\int_{0}^{1}\left(1-\tau^{2}\right)^{\frac{d-4}{2}}\Bigg(\frac{\bar{r}-r\tau}{\left(r^{2}+\bar{r}^{2}-2r\bar{r}\tau+(z-\bar{z})^{2}\right)^{\frac{d}{2}}}+\frac{\bar{r}+r\tau}{\left(r^{2}+\bar{r}^{2}+2r\bar{r}\tau+(z-\bar{z})^{2}\right)^{\frac{d}{2}}}\\ -\frac{\bar{r}-r\tau}{\left(r^{2}+\bar{r}^{2}-2r\bar{r}\tau+(z+\bar{z})^{2}\right)^{\frac{d}{2}}}-\frac{\bar{r}+r\tau}{\left(r^{2}+\bar{r}^{2}+2r\bar{r}\tau+(z+\bar{z})^{2}\right)^{\frac{d}{2}}}\Bigg)\mathop{}\!\mathrm{d}\tau.$ (6.5)
@@
$u_{r}(r,z)=-\frac{d-2}{2\pi}\int_{-\infty}^{\infty}\int_{0}^{\infty}\bar{r}^{d-2}(z-\bar{z})\omega(\bar{r},\bar{z})\int_{-1}^{1}\frac{\tilde{y}_{1}(1-\tilde{y}_{1}^{2})^{\frac{d-4}{2}}}{\left(r^{2}+\bar{r}^{2}-2r\bar{r}\tilde{y}_{1}+(z-\bar{z})^{2}\right)^{\frac{d}{2}}}\mathop{}\!\mathrm{d}\tilde{y}_{1}\mathop{}\!\mathrm{d}\bar{r}\mathop{}\!\mathrm{d}\bar{z}.$ (6.8)
@@
$u_{z}(r,z)=\frac{d-2}{2\pi}\int_{-\infty}^{\infty}\int_{0}^{\infty}\bar{r}^{d-2}\omega(\bar{r},\bar{z})\int_{-1}^{1}\frac{(r\tilde{y}_{1}-\bar{r})(1-\tilde{y}_{1}^{2})^{\frac{d-4}{2}}}{\left(r^{2}+\bar{r}^{2}-2r\bar{r}\tilde{y}_{1}+(z-\bar{z})^{2}\right)^{\frac{d}{2}}}\mathop{}\!\mathrm{d}\tilde{y}_{1}\mathop{}\!\mathrm{d}\bar{r}\mathop{}\!\mathrm{d}\bar{z}.$ (6.12)
@@
Suppose $u\in C\left([0,T_{max});H^{s}_{*}\left(\mathbb{R}^{d}\right)\right)\cap C^{1}\left([0,T_{max}),H^{s-1}_{*}\left(\mathbb{R}^{d}\right)\right),d\geq 3,s>2+\frac{d}{2}$, is a solution of the Euler equation. Then $\tilde{u}\in C\left([0,T_{max});H^{s}_{*}\left(\mathbb{R}^{d}\right)\right)\cap C^{1}\left([0,T_{max}),H^{s-1}_{*}\left(\mathbb{R}^{d}\right)\right)$ is also a solution of the Euler equation, where
@@
$\left(\partial_{t}\tilde{\omega}+\tilde{u}_{r}\partial_{r}\omega+\tilde{u}_{z}\partial_{z}\omega+\frac{k}{r}\tilde{u_{r}}\tilde{\omega}\right)(r,z,t)$ $=\left(-\partial_{t}\tilde{\omega}-\tilde{u}_{r}\partial_{r}\omega-\tilde{u}_{z}\partial_{z}\omega-\frac{k}{r}\tilde{u_{r}}\tilde{\omega}\right)(r,-z,t)$ (6.36)
$=0.$ (6.37)
@@
$\frac{\mathop{}\!\mathrm{d}}{\mathop{}\!\mathrm{d}t}\int_{0}^{\infty}\int_{0}^{\infty}\omega(r,z,t)\mathop{}\!\mathrm{d}r\mathop{}\!\mathrm{d}z$ $=\int_{0}^{\infty}\omega(0,z,t)u_{r}(0,z,t)\mathop{}\!\mathrm{d}z+\int_{0}^{\infty}\omega(r,0,t)u_{z}(r,0,t)\mathop{}\!\mathrm{d}r$ (6.43)
$=0,$ (6.44)
@@
$\frac{\mathop{}\!\mathrm{d}}{\mathop{}\!\mathrm{d}t}\int_{0}^{\infty}\int_{0}^{\infty}z\omega(r,z,t)\mathop{}\!\mathrm{d}r\mathop{}\!\mathrm{d}z=-\frac{1}{2}\int_{0}^{\infty}u_{z}(0,z)^{2}\mathop{}\!\mathrm{d}z-k\int_{0}^{\infty}\int_{0}^{\infty}\frac{1}{r}u_{r}(r,z)^{2}\mathop{}\!\mathrm{d}r\mathop{}\!\mathrm{d}z.$ (6.51)
@@
We will use the divergence form of the vorticity equation, $\partial_{t}\omega+\partial_{r}(\omega u_{r})+\partial_{z}(\omega u_{z})=0$. We can then compute that
@@
$\frac{\mathop{}\!\mathrm{d}}{\mathop{}\!\mathrm{d}t}\int_{0}^{\infty}\int_{0}^{\infty}r^{d-1}\omega(r,z,t)\mathop{}\!\mathrm{d}r\mathop{}\!\mathrm{d}z$ $=-\int_{0}^{\infty}\int_{0}^{\infty}r^{d-1}\left(\partial_{r}(\omega u_{r})+\partial_{z}(\omega u_{z})\right)$ (6.64)
$=(d-1)\int_{0}^{\infty}\int_{0}^{\infty}r^{d-2}u_{r}\omega\mathop{}\!\mathrm{d}r\mathop{}\!\mathrm{d}z$ (6.65)
$=(d-1)\int_{0}^{\infty}\int_{0}^{\infty}r^{d-2}u_{r}(\partial_{r}u_{z}-\partial_{z}u_{r})\mathop{}\!\mathrm{d}r\mathop{}\!\mathrm{d}z.$ (6.66)
@@
$\frac{\mathop{}\!\mathrm{d}}{\mathop{}\!\mathrm{d}t}\int_{0}^{\infty}\int_{0}^{\infty}r^{d-1}z\omega(r,z,t)\mathop{}\!\mathrm{d}r\mathop{}\!\mathrm{d}z$ $=-\int_{0}^{\infty}\int_{0}^{\infty}r^{d-1}z\left(\partial_{r}(\omega u_{r})+\partial_{z}(\omega u_{z})\right)\mathop{}\!\mathrm{d}r\mathop{}\!\mathrm{d}z$ (6.78)
$=\int_{0}^{\infty}\int_{0}^{\infty}\left((d-1)r^{d-2}z\omega u_{r}+r^{d-1}\omega u_{z}\right)\mathop{}\!\mathrm{d}r\mathop{}\!\mathrm{d}z.$ (6.79)

### v4

@@ 1 Introduction
As further evidence of potential singularity formation in very high dimensions, we prove that any solution of the Euler equation in $\mathbb{R}^{d}$ with a particular geometry—that is, axisymmetric, swirl-free, with a scalar vorticity $\omega^{0}(r,z)$ that is odd in $z$ and satisfying $\omega^{0}(r,z)\geq 0$ for all r,z>0r,z>0— blows up in finite time as long as a large enough proportion of energy remains in the vertical velocity component. Furthermore, the proportion of energy remaining in the vertical component necessary to guarantee finite-time blowup shrinks at a rate $\frac{1}{d}$, making the requirement very weak in high dimensions.
@@ 3.1 Preliminaries
$\nabla u(x)=\partial_{r}u_{r}(r,z)e_{r}\otimes e_{r}+\partial_{z}u_{z}(r,z)e_{z}\otimes e_{z}\\ +\partial_{r}u_{z}(r,z)e_{r}\otimes e_{z}+\partial_{z}u_{r}(r,z)e_{z}\otimes e_{r}+u_{r}(r,z)\nabla e_{r}.$ (3.5)
@@
$\Delta u=-\partial_{z}^{2}u_{r}e_{r}-\partial_{z}^{2}u_{z}e_{z}-\partial_{r}^{2}u_{z}e_{z}-\frac{k}{r}\partial_{r}u_{z}e_{z}-\operatorname{div}\left(\partial_{r}u_{r}e_{r}\otimes e_{r}\right)-\operatorname{div}\left(\frac{u_{r}}{r}\tilde{I}_{d}\right).$
@@
It is obvious that $\nabla\nabla\cdot(-\Delta)^{-1}v\in H^{s}_{as\&gr}$, because $\nabla\cdot v$ is an axisymmetric, scalar function, so it remains only to show that $2\operatorname{div}\nabla_{asym}(-\Delta)^{-1}v\in H^{s}_{*}$. Letting $A=\nabla_{asym}(-\Delta)^{-1}v$, we can see that
@@
Note that the above proposition shows that for all $v\in H^{s}_{as},P_{df}(v)\in H^{s}_{as}.$
@@ 3.2 Local existence of solutions
and then by proving strong convergence $u^{\epsilon}\to u$ in the space $C\left([0,T_{max}),H^{s}_{df}\left(\mathbb{R}^{d}\right)\right)\cap C^{1}\left([0,T_{max}),H^{s-1}_{df}\left(\mathbb{R}^{d}\right)\right)$. Note that $\mathcal{J}_{\epsilon}$ is a regularization operator given by convolution with a smooth, positive, compactly supported function of mass $1$. As long as $\mathcal{J}_{\epsilon}$ involves convolution with a radial function, then it will preserve the class of axisymmetric, swirl-free vector fields. The Picard iteration then guarantees that $u^{\epsilon}$ must be axisymmetric and swirl-free. Finally, $u$ must be axisymmetric and swirl-free, because it is trivial that this property is preserved by strong convergence in $H^{s}$.
@@ 4 The Biot-Savart law in higher dimensions
$\int_{\mathbb{S}^{d-2}}f(y_{1})\mathop{}\!\mathrm{d}S(y)$ $=m_{d-3}\int_{-1}^{1}\left(1-y_{1}^{2}\right)^{\frac{d-4}{2}}f(y_{1})\mathop{}\!\mathrm{d}y_{1}$ (4.2)
$=m_{d-3}\int_{0}^{\pi}f(\cos\theta)\sin^{d-3}\theta\mathop{}\!\mathrm{d}\theta,$ (4.3)
@@
$\int_{\mathbb{S}^{d-2}}f(y_{1})\mathop{}\!\mathrm{d}S(y)$ $=\int_{-1}^{1}\int_{|\tilde{y}|=\sqrt{1-y_{1}^{2}}}f(y_{1})\frac{1}{\sqrt{1-y_{1}^{2}}}\mathop{}\!\mathrm{d}S(\tilde{y})\mathop{}\!\mathrm{d}y_{1}$ (4.4)
$=\int_{-1}^{1}f(y_{1})\frac{1}{\sqrt{1-y_{1}^{2}}}\left|\mathbb{S}^{d-3}_{\sqrt{1-y_{1}^{2}}}\right|\mathop{}\!\mathrm{d}y_{1}.$ (4.5)
@@
$u_{r}(r,z)=-\frac{d-2}{2\pi}\int_{-\infty}^{\infty}\int_{0}^{\infty}\bar{r}^{d-2}(z-\bar{z})\omega(\bar{r},\bar{z})\\ \int_{-1}^{1}\frac{y_{1}(1-y_{1}^{2})^{\frac{d-4}{2}}}{\left(r^{2}+\bar{r}^{2}-2r\bar{r}y_{1}+(z-\bar{z})^{2}\right)^{\frac{d}{2}}}\mathop{}\!\mathrm{d}y_{1}\mathop{}\!\mathrm{d}\bar{r}\mathop{}\!\mathrm{d}\bar{z}$ (4.6)
@@
$u_{r}(r,z)=-\frac{d-2}{2\pi}\int_{-\infty}^{\infty}\int_{0}^{\infty}\bar{r}^{d-2}(z-\bar{z})\omega(\bar{r},\bar{z})\\ \int_{0}^{\pi}\frac{\sin^{d-3}\theta\cos\theta}{\left(r^{2}+\bar{r}^{2}-2r\bar{r}\cos\theta+(z-\bar{z})^{2}\right)^{\frac{d}{2}}}\mathop{}\!\mathrm{d}\theta\mathop{}\!\mathrm{d}\bar{r}\mathop{}\!\mathrm{d}\bar{z}$ (4.7)
@@
$u_{z}(r,z)=\frac{d-2}{2\pi}\int_{-\infty}^{\infty}\int_{0}^{\infty}\bar{r}^{d-2}\omega(\bar{r},\bar{z})\\ \int_{0}^{\pi}\frac{\sin^{d-3}\theta(r\cos\theta-\bar{r})}{\left(r^{2}+\bar{r}^{2}-2r\bar{r}\cos\theta+(z-\bar{z})^{2}\right)^{\frac{d}{2}}}\mathop{}\!\mathrm{d}\theta\mathop{}\!\mathrm{d}\bar{r}\mathop{}\!\mathrm{d}\bar{z}.$ (4.9)
@@
$u_{r}(r,z)=-\alpha_{d}\int_{-\infty}^{\infty}\int_{0}^{\infty}\bar{r}^{d-2}(z-\bar{z})\omega(\bar{r},\bar{z})\\ \int_{\mathbb{S}^{d-2}}\frac{y_{1}}{\left(r^{2}+\bar{r}^{2}-2r\bar{r}y_{1}+(z-\bar{z})^{2}\right)^{\frac{d}{2}}}\mathop{}\!\mathrm{d}S(y)\mathop{}\!\mathrm{d}\bar{r}\mathop{}\!\mathrm{d}\bar{z}.$ (4.10)
@@ 5 Conditional regularity in four and higher dimensions
$\|\omega\circ X(\cdot,t)\|_{L^{q}\left(\mathbb{R}^{d}\right)}^{q}$ $=\|\omega\circ X(\cdot,t)\|_{L^{q}\left(\mathcal{C}_{R}\right)}^{q}+\|\omega\circ X(\cdot,t)\|_{L^{q}\left(\mathcal{C}_{R}^{c}\right)}^{q}$
$\leq\left(\left\|\frac{\omega^{0}}{r^{k}}\right\|_{L^{q}\left(\mathcal{C}_{R}\right)}^{q}R^{kq}+\left\|\omega^{0}\right\|_{L^{q}\left(\mathcal{C}_{R}^{c}\right)}^{q}\right)\left(1+\frac{1}{R}\int_{0}^{t}\|u_{r}^{+}(\cdot,\tau)\|_{L^{\infty}}\mathop{}\!\mathrm{d}\tau\right)^{kq}$
@@
$\|u_{r}^{+}(\cdot,t)\|_{L^{\infty}}\leq C_{d}\left\|\frac{\omega^{0}}{r^{k}}\right\|_{L^{\infty}}^{\frac{1}{2}}\left(\left\|\omega^{0}\right\|_{L^{1}(\mathcal{C}_{R}^{c})}+R^{k}\left\|\frac{\omega^{0}}{r^{k}}\right\|_{L^{1}(\mathcal{C}_{R})}\right)^{\frac{1}{2}}\\ \left(1+\frac{1}{R}\int_{0}^{t}\|u_{r}^{+}(\cdot,\tau)\|_{L^{\infty}}\mathop{}\!\mathrm{d}\tau\right)^{\frac{k}{2}}.$ (5.21)
@@
$\|u_{r}^{+}(\cdot,t)\|_{L^{\infty}}$ $\leq C_{d}\left\|\frac{\omega(\cdot,t)}{r^{k}}\right\|_{L^{\infty}}^{\frac{1}{2}}\|\omega(\cdot,t)\|_{L^{1}}^{\frac{1}{2}}$
$\leq\left\|\frac{\omega^{0}}{r^{k}}\right\|_{L^{\infty}}^{\frac{1}{2}}\left(\left\|\omega^{0}\right\|_{L^{1}(\mathcal{C}_{R}^{c})}+R^{k}\left\|\frac{\omega^{0}}{r^{k}}\right\|_{L^{1}(\mathcal{C}_{R})}\right)^{\frac{1}{2}}\left(1+\frac{1}{R}\int_{0}^{t}\|u_{r}^{+}(\cdot,\tau)\|_{L^{\infty}}\mathop{}\!\mathrm{d}\tau\right)^{\frac{k}{2}}.$
@@ 6 Anti-parallel vortex tubes
$H(r,z,\bar{r},\bar{z})=\int_{0}^{1}\Bigg(\frac{z+\bar{z}}{\left(r^{2}+\bar{r}^{2}-2r\bar{r}\tau+(z+\bar{z})^{2}\right)^{\frac{d}{2}}}+\frac{z-\bar{z}}{\left(r^{2}+\bar{r}^{2}+2r\bar{r}\tau+(z-\bar{z})^{2}\right)^{\frac{d}{2}}}\\ -\frac{z+\bar{z}}{\left(r^{2}+\bar{r}^{2}+2r\bar{r}\tau+(z+\bar{z})^{2}\right)^{\frac{d}{2}}}-\frac{z-\bar{z}}{\left(r^{2}+\bar{r}^{2}-2r\bar{r}\tau+(z-\bar{z})^{2}\right)^{\frac{d}{2}}}\Bigg)\tau\left(1-\tau^{2}\right)^{\frac{d-4}{2}}\mathop{}\!\mathrm{d}\tau,$ (6.3)
@@
$G(r,z,\bar{r},\bar{z})=\int_{0}^{1}\Bigg(\frac{\bar{r}-r\tau}{\left(r^{2}+\bar{r}^{2}-2r\bar{r}\tau+(z-\bar{z})^{2}\right)^{\frac{d}{2}}}+\frac{\bar{r}+r\tau}{\left(r^{2}+\bar{r}^{2}+2r\bar{r}\tau+(z-\bar{z})^{2}\right)^{\frac{d}{2}}}\\ -\frac{\bar{r}-r\tau}{\left(r^{2}+\bar{r}^{2}-2r\bar{r}\tau+(z+\bar{z})^{2}\right)^{\frac{d}{2}}}-\frac{\bar{r}+r\tau}{\left(r^{2}+\bar{r}^{2}+2r\bar{r}\tau+(z+\bar{z})^{2}\right)^{\frac{d}{2}}}\Bigg)\left(1-\tau^{2}\right)^{\frac{d-4}{2}}\mathop{}\!\mathrm{d}\tau.$ (6.5)
@@
$u_{r}(r,z)=-\frac{d-2}{2\pi}\int_{-\infty}^{\infty}\int_{0}^{\infty}\bar{r}^{d-2}(z-\bar{z})\omega(\bar{r},\bar{z})\\ \int_{-1}^{1}\frac{\tilde{y}_{1}(1-\tilde{y}_{1}^{2})^{\frac{d-4}{2}}}{\left(r^{2}+\bar{r}^{2}-2r\bar{r}\tilde{y}_{1}+(z-\bar{z})^{2}\right)^{\frac{d}{2}}}\mathop{}\!\mathrm{d}\tilde{y}_{1}\mathop{}\!\mathrm{d}\bar{r}\mathop{}\!\mathrm{d}\bar{z}.$ (6.8)
@@
$u_{z}(r,z)=\frac{d-2}{2\pi}\int_{-\infty}^{\infty}\int_{0}^{\infty}\bar{r}^{d-2}\omega(\bar{r},\bar{z})\int_{-1}^{1}\frac{(r\tilde{y}_{1}-\bar{r})(1-\tilde{y}_{1}^{2})^{\frac{d-4}{2}}}{\left(r^{2}+\bar{r}^{2}-2r\bar{r}\tilde{y}_{1}+(z-\bar{z})^{2}\right)^{\frac{d}{2}}}\mathop{}\!\mathrm{d}\tilde{y}_{1}\mathop{}\!\mathrm{d}\bar{r}\mathop{}\!\mathrm{d}\bar{z}.$
@@
Suppose $u\in C\left([0,T_{max});H^{s}_{*}\left(\mathbb{R}^{d}\right)\right)\cap C^{1}\left([0,T_{max}),H^{s-1}_{*}\left(\mathbb{R}^{d}\right)\right),d\geq 3,s>2+\frac{d}{2}$, is a solution of the Euler equation. Then the velocity $\tilde{u}\in C\left([0,T_{max});H^{s}_{*}\left(\mathbb{R}^{d}\right)\right)\cap C^{1}\left([0,T_{max}),H^{s-1}_{*}\left(\mathbb{R}^{d}\right)\right)$ is also a solution of the Euler equation, where
@@
$\left(\partial_{t}\tilde{\omega}+\tilde{u}_{r}\partial_{r}\omega+\tilde{u}_{z}\partial_{z}\omega+\frac{k}{r}\tilde{u_{r}}\tilde{\omega}\right)(r,z,t)$ $=\left(-\partial_{t}\tilde{\omega}-\tilde{u}_{r}\partial_{r}\omega-\tilde{u}_{z}\partial_{z}\omega-\frac{k}{r}\tilde{u_{r}}\tilde{\omega}\right)(r,-z,t)$
$=0.$
@@
$\frac{\mathop{}\!\mathrm{d}}{\mathop{}\!\mathrm{d}t}\int_{0}^{\infty}\int_{0}^{\infty}\omega(r,z,t)\mathop{}\!\mathrm{d}r\mathop{}\!\mathrm{d}z$ $=\int_{0}^{\infty}\omega(0,z,t)u_{r}(0,z,t)\mathop{}\!\mathrm{d}z+\int_{0}^{\infty}\omega(r,0,t)u_{z}(r,0,t)\mathop{}\!\mathrm{d}r$
$=0,$
@@
$\frac{\mathop{}\!\mathrm{d}}{\mathop{}\!\mathrm{d}t}\int_{0}^{\infty}\int_{0}^{\infty}z\omega(r,z,t)\mathop{}\!\mathrm{d}r\mathop{}\!\mathrm{d}z=-\frac{1}{2}\int_{0}^{\infty}u_{z}(0,z)^{2}\mathop{}\!\mathrm{d}z-k\int_{0}^{\infty}\int_{0}^{\infty}\frac{1}{r}u_{r}(r,z)^{2}\mathop{}\!\mathrm{d}r\mathop{}\!\mathrm{d}z.$
@@
We will use the divergence form of the vorticity equation,
We can then compute that
@@
$\frac{\mathop{}\!\mathrm{d}}{\mathop{}\!\mathrm{d}t}\int_{0}^{\infty}\int_{0}^{\infty}r^{d-1}\omega(r,z,t)\mathop{}\!\mathrm{d}r\mathop{}\!\mathrm{d}z$ $=-\int_{0}^{\infty}\int_{0}^{\infty}r^{d-1}\left(\partial_{r}(\omega u_{r})+\partial_{z}(\omega u_{z})\right)$
$=(d-1)\int_{0}^{\infty}\int_{0}^{\infty}r^{d-2}u_{r}\omega\mathop{}\!\mathrm{d}r\mathop{}\!\mathrm{d}z$
$=(d-1)\int_{0}^{\infty}\int_{0}^{\infty}r^{d-2}u_{r}(\partial_{r}u_{z}-\partial_{z}u_{r})\mathop{}\!\mathrm{d}r\mathop{}\!\mathrm{d}z.$
@@
$\frac{\mathop{}\!\mathrm{d}}{\mathop{}\!\mathrm{d}t}\int_{0}^{\infty}\int_{0}^{\infty}r^{d-1}z\omega(r,z,t)\mathop{}\!\mathrm{d}r\mathop{}\!\mathrm{d}z$ $=-\int_{0}^{\infty}\int_{0}^{\infty}r^{d-1}z\left(\partial_{r}(\omega u_{r})+\partial_{z}(\omega u_{z})\right)\mathop{}\!\mathrm{d}r\mathop{}\!\mathrm{d}z$
$=\int_{0}^{\infty}\int_{0}^{\infty}\left((d-1)r^{d-2}z\omega u_{r}+r^{d-1}\omega u_{z}\right)\mathop{}\!\mathrm{d}r\mathop{}\!\mathrm{d}z.$